OpenStudy (anonymous):

Write the equation of the line that passes through the points (–1, –4) and (2, 5).

OpenStudy (anonymous):

m = 3 y - 5= 3(x + 2) y = 3x + 11

OpenStudy (amistre64):



OpenStudy (amistre64):

y - 5= 3(x + 2) ^^ needs to be subtract 2 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

typos happen, but its good to know when it does :)